The sum of the years of the mother and the daughter is 56. 5 years ago the mother was 2 times older than the daughter. How old are they now?
The solving looks something like
x+y=56
(x-5)=(y-5)*2
Please explain thoroughly and step by step :)

Answer by Alan3354(69443)

The sum of the years of the mother and the daughter is 56. 5 years ago the mother was 2 times older than the daughter. How old are they now?
The solving looks something like
x+y=56
(x-5)=(y-5)*2
---------------
1 time older = 100% older = 2x as old
2 times older = 200% older = 3x as old
==========
I would use m & d instead of x & y
----
m + d = 56
m-5 = 3*(d-5) = 3d - 15 --> m - 3d = -10
---
m - 3d = -10
m + d = 56
--------------------- Subtract
-4d = -66
d = 16.5
m = 39.5